OpenStudy (anonymous):

.4% of 36 equals ? and how please break it down

OpenStudy (curry):

52 / 5 = 10.4

OpenStudy (anonymous):

.4/100 = x/36

OpenStudy (anonymous):

cross multiply and solve for x

OpenStudy (anonymous):

oops. .4%=.004 so you need \[.004\times 36\]

OpenStudy (dumbcow):

usually "of" implies multiplication .4% = .004 (just move decimal 2 to the left) -> (.004)*36 = x

OpenStudy (anonymous):

36*.4=100x

OpenStudy (anonymous):

14.4=100x

OpenStudy (anonymous):

.144=x

OpenStudy (anonymous):

i get .114

OpenStudy (anonymous):

so 14 cents or 15 cents, depending on how u round

OpenStudy (anonymous):

what is the ratio of 8:5 = %? and how please

OpenStudy (anonymous):

\[\frac{8}{5}=\frac{x}{100}\]

OpenStudy (anonymous):

cross multiply again

OpenStudy (anonymous):

800=5x

OpenStudy (anonymous):

160=x

OpenStudy (anonymous):

8 is 160% of 5